OpenStudy (anonymous):

Hey guys! help me ;; given that: g(g(x))=6x-8 how do we find g(x)? Thanks guys love you all

OpenStudy (anonymous):

I have the answer... \[g(x)=\sqrt{6}x-\frac{ 8 }{ \sqrt{6}+1 }\] but I have no idea how to get it

OpenStudy (freckles):

hey

OpenStudy (freckles):

I will try to think with ya

OpenStudy (freckles):

hmm.. is this calculus or algebra?

OpenStudy (freckles):

I guess we can try something since we have g(g(x)) is linear than g(x) is linear I believe so we know g(x)=ax+b

OpenStudy (freckles):

\[g(x)=ax+b \\ g(g(x))=a(ax+b)+b \\ a^2x+ab+b=6x-8\]

OpenStudy (freckles):

so this mean comparing both sides we have: \[a^2=6 \\ \text{ and } \\ ab+b=-8\]

OpenStudy (freckles):

find a then you can find b

OpenStudy (anonymous):

This is pre-calc, so algebra Thank you so much freckles!

OpenStudy (freckles):

Np